A spotlight is 5 feet from one wall of a room and 10 feet from the other wall at right angles to it. How many feet is it from the intersection of the two walls?
A)15
B)7.07
C)11.18
D)14.14
E)22.36

Answer by stanbon(75887)

A spotlight is 5 feet from one wall of a room and 10 feet from the other wall at right angles to it. How many feet is it from the intersection of the two walls?
----
Let the walls be the x and y axis of a coordinate system.
Plot a point at (5,10)
---
Find the distance of the point from the origin.
---
d^2 = 5^2 + 10^2
d = sqrt(125)
d = 5*sqrt(5)
d = 11.18
=====================
Cheers,
Stan H.
---------------------
A)15
B)7.07
C)11.18
D)14.14
E)22.36